OpenStudy (sarahc):

HELPPPPPP The graph of a limacon curve is given. Without using your graphing calculator, determine which equation is correct for the graph. [-5, 5] by [-5, 5] r = 2 - 2 sin θ r = 3 - 2 sin θ r = 2 + 2 sin θ r = 1 - 3 sin θ

OpenStudy (sarahc):

I got a?

OpenStudy (sarahc):

@Directrix

Directrix (directrix):

Where is the graph: The graph of a limacon curve is given. Please post it. Thanks.

OpenStudy (sarahc):

@Directrix Sorry!!

Directrix (directrix):

I'll see if I get A.

Directrix (directrix):

i'm going with Option A @sarahc

OpenStudy (kmeis002):

Since a limacon follows the following forms: \[r = a\pm b \cos(\theta) \quad \text{or} \quad r = a \pm b \sin(\theta) \] We know it must be a function of sine since its largest radius is at \[ \theta = \frac{3\pi}{2} \] Since sin(x) is minimized at that angle, this would return the smallest radius. So it must be \[r = a - b \sin(\theta) \] As for a and b, we can see from the graph that the radius does become 0 at one point (r=0). It is also never negative, because this creates the second inner loop in a limacon. Since the range of sin(x) is [-1,1], the only way for the radius to be greater than or equal to zero is if: \[ a = b\] This gives us A as the answer

OpenStudy (anonymous):

it does not use sine. this is from the textbook

OpenStudy (anonymous):

r = 3 - cos θ 100%